Refer to the following selected financial information from McCormik, LLC. Compute the companys current ratio for Year 2. CasRefer to the following selected financial information from McCormik, LLC. Compute the companys inventory turnover for Year 2Refer to the following selected financial information from McCormik, LLC. Compute the companys acid-test ratio for Year 2. C
Answer #1

1) Current ratio = Current assets/Current liabilities = 258500/112400 = 2.30:1

So answer is a) 2.30

2) Inventory turnover = Cost of goods sold/average inventory = 394000/127000 = 3.10

So answer is a) 3.10

3) Acid test ratio = Quick assets/current liabilities = 238600/105400 = 2.26

So answer is c) 2.26